OpenStudy (anonymous):

Radicals A rectangular garden is 16 feet long. If you walk diagonally across the garden, you would walk 20 feet. How many feet wide is the garden?

OpenStudy (oaktree):

Essentially, we have a case of a right triangle with two given sides, which can be solved trivially by means of the Pythagorean Theorem, or \[a^2+b^2=c^2\]We know that b=16 and c=20. Hence, we have \[a=\sqrt{c^2-b^2} = \sqrt{20^2-16^2} = \sqrt{144} = 12.\]So the width of the garden is 12 feet.